Two people set off to make the same journey between two villages 5 km apart. One goes by car and the other by cycle.
1. The cyclist cycles at a steady speed of 5m/s. How many metres does the cyclist travel? How many seconds does it take the cyclist to complete the journey?
2. The car sets off 500 seconds after the cyclist and travels at a constant speed of 20/s. How many seconds does it take the far to complete the journey?

Answer by mananth(16946)

1. The cyclist cycles at a steady speed of 5m/s. How many metres does the cyclist travel? How many seconds does it take the cyclist to complete the journey?
distance is 5 km
1 km = 1000 m
5 km = 5000 m
time taken by cyclist = d/r => 5000/5 - 1000 seconds
2. The car sets off 500 seconds after the cyclist and travels at a constant speed of 20/s. How many seconds does it take the far to complete the journey?
distance = 5000 m
rate = 20 m/s
time = 5000/20 = 100 seconds